31 October 2007

Short update.

His temperature has been kept below 38.4º C today. Admittedly with the help of paracetamol, but still. CRP slightly up again, bot nothing dramatic. The day started pretty much as it ended yesterday, i.e. with a couple of helpings of proper nose bleeding.

The doctors think Abbe’s temperature is a totally natural reaction to the newly transplated homograft (a human vessel from a donator) so yesterday they started a course of ASA (aspirin) in order to deal with it. The problem is, aspirin also thins the blood which makes the nosebleeding even worse. So, today they stopped the ASA again. Catch 22 or what.

By the way, I have been meaning for days to try and explain what they have done to Abbe this time. The actual operation, I mean. It’ll come, I promise. Maybe tomorrow.

No comments: